OpenStudy (anonymous):

rewrite the following fraction by rationalizing the denominator. 7/(√3x + √(x-4))

OpenStudy (anonymous):

multiply the denominator by itself and also the numerator by the dominator

OpenStudy (anonymous):

this is what i had but my teacher said i have to simplify it more 7(√3x-√(x-4) / 2(x+2)

OpenStudy (anonymous):

I know tat u have to do that

OpenStudy (anonymous):

start by distributing the 7

OpenStudy (anonymous):

ok I distributed it to to the numerator

OpenStudy (anonymous):

somebody help me

OpenStudy (austinl):

\[\frac{7}{\sqrt{3x}+\sqrt{x-4}}\]or? \[\frac{7}{\sqrt{3}x+\sqrt{x-4}}\]

OpenStudy (anonymous):

1st one

OpenStudy (austinl):

\[\frac{7\sqrt{3}\sqrt{x}-7\sqrt{x-4}}{2x+4}\] This is your final answer, which you get by multiplying both numerator and denominator by the denominator.

OpenStudy (anonymous):

so u would put both the 3 and x under two different radical signs??

OpenStudy (austinl):

You can, it doesn't really matter too much.

OpenStudy (anonymous):

so if I write it JUST like you did, it would be correct??

OpenStudy (austinl):

Technically, yes.

OpenStudy (anonymous):

k